# Bootstrap3D: Improving 3D Content Creation with Synthetic Data

### v1

@@ B.1 Caption Analysis
Fig. 8 and  9 provide a visualization of the root noun-verb pairs for the captions generated by GPT-4V [55] and MV-LLaVA. It’s clear to see that the diversity and linguistic expression of the captions produced by MV-LLaVA are highly matched with those of GPT-4V. We believe the highly detailed description focusing on object’s texture, shape and color have potential usage beyond training multi-view diffusion model in the field like object texturing [26] and stylization [65] in Computer Graphics. MV-LLaVA can also serve as free and efficient 3D object assistant comparable with GPT-4V for future research of 3D content creation.
Figure 10: Histogram Visualization of the Caption Length compared with Cap3D [52]
@@
Fig. 10 visualizes the histogram of caption length compared with Cap3D [52]. We fine-tune MV-LLaVA to generate two different lengths suitable for different diffusion architecture, namely CLIP-based text encoding [6, 58] with 77 token length and T5 based text encoding [12, 11] with 120 token length. Both excel the length of Cap3D with less hallucinations.
B.2 Estimated Quality Analysis
@@
We count the number of multi-view images from different data sources, namely 660K from Objaverse, 500K from SV3D [76] and 500K from Zero123++ [66] generated by our Bootstrap3D pipeline. Result are shown in Fig.11. For Objaverse and SV3D, the assigned score is normal and we use score 4 and score 5 multi-view images as high quality data for training. However, for Zero123++, most objects are assigned with score greater than 3. We attribute this phenomenon to the fact that Zero123++ tend to generate objects with less motion blurring but more stretching and deformation compared to SV3D. Joint training of MV-LLaVA on three different data source lead to higher and more focused distribution for Zero123++’s multi-view images. For this part of synthetic data, we leave only score 5 multi-view images as high quality data.
Appendix C Quality of MV-LLaVA
C.1 Choice of number of unfrozen layers of vision encoder.
@@ Appendix C Quality of MV-LLaVA
Figure 12: Qualitative results of unfreeze final layers of CLIP [61] vision encoder compared to original fixed vision encoder setting in LLaVA [46].
@@
Inspired by ShareGPT-4V [13], we unfreeze selected final layers of the CLIP [46] vision encoder during the initial phase of vision language alignment. The CLIP-L/14 model used for LLaVA [46] contains 24 transformer layers. We selectively unfreeze some of final layers to enable the CLIP model to focus more on details such as texture of multi-view images. After qualitative manual screening, we select to unfreeze eight layers to yield better results. Fig. 12 illustrates the differences between unfreezing eight layers and not unfreezing any (the original training setting of LLaVA [46]). The red sections highlight the erroneous hallucinations occurring when the vision encoder remains fully unchanged, while the green sections indicate accurate descriptions of the image content. This demonstrates that partially unfreezing the vision encoder can produce more precise captions and reduce some hallucinations.
C.2 Quantitative quality study
@@
To test the quality of our MV-LLaVA. We propose two quantitative study over the quality of captions and the alignment of quality estimation with human experts. In first study, we randomly picked 200 object from Objaverse [20] and exclude training data of MV-LLaVA. We use GPT4-V [55] and MV-LLaVA to generate descriptive captions for each object. We invite human volunteers to choose their preference over shuffled captions. Results are shown in Tab. 6, where MV-LLaVA shows comparable captioning ability with powerful GPT4-V [55], which is essential to generate millions of high quality image-text pairs for the training of text to multi-view image diffusion model.
@@
Second experiment studies MV-LLaVA’s ability in quality estimation of both 3D assets and generated multi-view images. We invite human volunteers to estimate the quality of multi-view images rendered from Objaverse [20] or generated by SV3D [76]. As there is no golden standard for multi quality classification, We ask them to separate the randomly select multi-view images into approximately two half and serve as GT quality. We use MV-LLaVA to estimate the quality of these images and generate confusion matrix. Results are shown in Tab.7. Given the great amount of source data of 3D assets and infinite synthetic data, we care more about the false positive rate, as these data will be mixed into training data. In this observation, we highlight the false positive rate of over 20% for SV3D [76] generated multi-view images. This result align with the observation of inevitable motion blurring of SV3D [76]. To leverage this part of data source for data diversity without hurting the final quality. We propose Training Noise Reschedule to avoid samplings from these synthetic data when time step is small.
@@
Table 6: Human evaluation on the quality of generated captions from MV-LLaVA vs. GPT4-Vision [55] over 200 validation samples from Objaverse [20].
Preference GPT4-Vision [55] MV-LLaVA Comparable
@@
C.3 Qualitative caption quality study
We selective compare some of the captions generated by Cap3D [52] and MV-LLaVA in Fig. 13. Our MV-LLaVA can generate more detailed descriptive captions with less hallucinations.
Figure 13: Caption comparison with Cap3D [52]. Our MV-LLaVA can generate long captions that faithfully describing 3D assets from different perspectives like color, geometry and texture.
Figure 14: Randomly picked multi-view images with different scores from 500k synthetic data generated by SV3D [76].
Figure 15: Randomly picked multi-view images with different scores from 500k synthetic data generated by Zero123++ [66].
Figure 16: Randomly picked multi-view images with different scores from 660k Objaverse [20] 3D assets.
Appendix D Details of Prompt Design
D.1 Prompts for GPT-4V for Quality Check
Figure 17: Prompt for GPT-4V to generate caption and estimate quality of multi-view images from SV3D [76], zero123++ [66] and Objaverse [20].
@@ D.1 Prompts for GPT-4V for Quality Check
D.2 Prompts for MV-LLaVA Instruct Tuning
Table 8: Instruct tuning prompt for SV3D [76] and Zero123++ [66] multi-view images
@@ D.2 Prompts for MV-LLaVA Instruct Tuning
quality estimation What do you think about the overall quality of view consistency of three synthesized novel views? Choosing from "poor", "relatively poor", "boardline", "relatively good", "good", "perfect".
Table 9: Instruct tuning prompt for Objaverse [20] rendered multi-view images
@@
quality estimation What do you think about the overall quality of this 3D model? Choosing from "poor", "relatively poor", "boardline", "relatively good", "good", "perfect".
scale tag What do you think about the scale of the 3D model represents? Choosing from "single_object", "multi-object", "small_scene", "large_scene".
style tag What do you think about the overall style of the 3D model? Choosing from "CAD", "Cartoon", "Photo_realistic".
Figure 18: A test conversation with GPT-4V [55] of evaluating generated objects.
Appendix E GPT-4V based 3D Object Generation Evaluation.
We adopt method proposed in GPTeval3D [85] for more thorough and human-aligned evaluation of the quality of generated object by different methods. A full test case is shown in Fig. 18. Left 9-view image is rendered from object generated by Bootstrap3D and the right one generated by Instant3D [41]. We ask GPT-4V to mainly evaluate through comparison based on three dimensions: text-image alignment, low-level texture quality and 3D plausibility. The answer of GPT-4V shows its in depth perception ability of given reasonable comparison well aligned with human preference. We thus choose to use GPT-4V rather than human volunteers to give reasonable evaluation.
We adopt the 110 test prompts proposed in GPTeval3D [85] to test Bootstrap3D generated object comparing with Instant3D [41], Zero123++ [66] and MVDream [67]. For each methods, we conditioned model based on 110 test prompts with 4 different seeds, with each methods generates 440 objects, we make 1-to-1 comparison following aforementioned test setting. Results are reported in Tab. 10. Except MVDream [67] (SDS) (which generates single object consuming 30 mins while Bootstrap3D only need 5 seconds.). Bootstrap3D excels in all three evaluation dimensions, which proves the ability of Bootstrap3D in creating high quality 3D objects.
Table 10: GPT-4V based evaluation result. the result is in format of "number of objects preferred geneated by Bootstrap3D/ that of other methods". Cases when GPT cannot answer the question or generates "cannot decide" answer are excluded.
@@
Compared to Instant3D [41] (unofficial) 247 / 116 202 / 162 259 / 110
Compared to Zero123++ [66] 192 / 143 210 / 161 231 / 139
Compared to MVDream [67] (GRM) 290 / 71 245 / 131 284 / 102
Compared to MVDream [67] (SDS) 188 / 155 173 / 190 192 / 150
Appendix F More Results Visualization
F.1 Comparison with Other Methods
@@
F.2 Visualization of Generated objects with Different Styles

### v2

@@ A.3.1 Caption Analysis
Fig. 10 and  11 provide a visualization of the root noun-verb pairs for the captions generated by GPT-4V (OpenAI, 2023a) and MV-LLaVA. It’s clear to see that the diversity and linguistic expression of the captions produced by MV-LLaVA are highly matched with those of GPT-4V. We believe the highly detailed description focusing on object’s texture, shape and color have potential usage beyond training multi-view diffusion model in the field like object texturing (Fang et al., 2024) and stylization (Sharma et al., 2023) in Computer Graphics. MV-LLaVA can also serve as free and efficient 3D object assistant comparable with GPT-4V for future research of 3D content creation.
Figure 12: Histogram Visualization of the Caption Length compared with Cap3D (Luo et al., 2024)
@@
Fig. 12 visualizes the histogram of caption length compared with Cap3D (Luo et al., 2024). We fine-tune MV-LLaVA to generate two different lengths suitable for different diffusion architecture, namely CLIP-based text encoding (Blattmann et al., 2023; Podell et al., 2023) with 77 token length and T5 based text encoding (Chen et al., 2023a; 2024a) with 120 token length. Both excel the length of Cap3D with less hallucinations.
A.3.2 Estimated Quality Analysis
@@
We count the number of multi-view images from different data sources, namely 660K from Objaverse, 500K from SV3D (Voleti et al., 2024) and 500K from Zero123++ (Shi et al., 2023a) generated by our Bootstrap3D pipeline. Result are shown in Fig.13. For Objaverse and SV3D, the assigned score is normal and we use score 4 and score 5 multi-view images as high quality data for training. However, for Zero123++, most objects are assigned with score greater than 3. We attribute this phenomenon to the fact that Zero123++ tend to generate objects with less motion blurring but more stretching and deformation compared to SV3D. Joint training of MV-LLaVA on three different data source lead to higher and more focused distribution for Zero123++’s multi-view images. For this part of synthetic data, we leave only score 5 multi-view images as high quality data.
A.4 Quality of MV-LLaVA
A.4.1 Choice of number of unfrozen layers of vision encoder.
@@ A.4 Quality of MV-LLaVA
Figure 14: Qualitative results of unfreeze final layers of CLIP (Radford et al., 2021) vision encoder compared to original fixed vision encoder setting in LLaVA (Liu et al., 2024a).
@@
Inspired by ShareGPT-4V (Chen et al., 2023b), we unfreeze selected final layers of the CLIP (Liu et al., 2024a) vision encoder during the initial phase of vision language alignment. The CLIP-L/14 model used for LLaVA (Liu et al., 2024a) contains 24 transformer layers. We selectively unfreeze some of final layers to enable the CLIP model to focus more on details such as texture of multi-view images. After qualitative manual screening, we select to unfreeze eight layers to yield better results. Fig. 14 illustrates the differences between unfreezing eight layers and not unfreezing any (the original training setting of LLaVA (Liu et al., 2024a)). The red sections highlight the erroneous hallucinations occurring when the vision encoder remains fully unchanged, while the green sections indicate accurate descriptions of the image content. This demonstrates that partially unfreezing the vision encoder can produce more precise captions and reduce some hallucinations.
A.4.2 Quantitative quality study
@@
To test the quality of our MV-LLaVA. We propose two quantitative study over the quality of captions and the alignment of quality estimation with human experts. In first study, we randomly picked 200 object from Objaverse (Deitke et al., 2023) and exclude training data of MV-LLaVA. We use GPT4-V (OpenAI, 2023a) and MV-LLaVA to generate descriptive captions for each object. We invite human volunteers to choose their preference over shuffled captions. Results are shown in Tab. 6, where MV-LLaVA shows comparable captioning ability with powerful GPT4-V (OpenAI, 2023a), which is essential to generate millions of high quality image-text pairs for the training of text to multi-view image diffusion model.
@@
Second experiment studies MV-LLaVA’s ability in quality estimation of both 3D assets and generated multi-view images. We invite human volunteers to estimate the quality of multi-view images rendered from Objaverse (Deitke et al., 2023) or generated by SV3D (Voleti et al., 2024). As there is no golden standard for multi quality classification, We ask them to separate the randomly select multi-view images into approximately two half and serve as GT quality. We use MV-LLaVA to estimate the quality of these images and generate confusion matrix. Results are shown in Tab.7. Given the great amount of source data of 3D assets and infinite synthetic data, we care more about the false positive rate, as these data will be mixed into training data. In this observation, we highlight the false positive rate of over 20% for SV3D (Voleti et al., 2024) generated multi-view images. This result align with the observation of inevitable motion blurring of SV3D (Voleti et al., 2024). To leverage this part of data source for data diversity without hurting the final quality. We propose Training Noise Reschedule to avoid samplings from these synthetic data when time step is small.
@@
Table 6: Human evaluation on the quality of generated captions from MV-LLaVA vs. GPT4-Vision (OpenAI, 2023a) over 200 validation samples from Objaverse (Deitke et al., 2023).
Preference GPT4-Vision (OpenAI, 2023a) MV-LLaVA Comparable
@@
A.4.3 Qualitative caption quality study
We selective compare some of the captions generated by Cap3D (Luo et al., 2024) and MV-LLaVA in Fig. 15. Our MV-LLaVA can generate more detailed descriptive captions with less hallucinations.
Figure 15: Caption comparison with Cap3D (Luo et al., 2024). Our MV-LLaVA can generate long captions that faithfully describing 3D assets from different perspectives like color, geometry and texture.
Figure 16: Randomly picked multi-view images with different scores from 500k synthetic data generated by SV3D (Voleti et al., 2024).
Figure 17: Randomly picked multi-view images with different scores from 500k synthetic data generated by Zero123++ (Shi et al., 2023a).
Figure 18: Randomly picked multi-view images with different scores from 660k Objaverse (Deitke et al., 2023) 3D assets.
A.5 Details of Prompt Design
A.5.1 Prompts for GPT-4V for Quality Check
Figure 19: Prompt for GPT-4V to generate caption and estimate quality of multi-view images from SV3D (Voleti et al., 2024), zero123++ (Shi et al., 2023a) and Objaverse (Deitke et al., 2023).
@@ A.5.1 Prompts for GPT-4V for Quality Check
A.5.2 Prompts for MV-LLaVA Instruct Tuning
Table 8: Instruct tuning prompt for SV3D (Voleti et al., 2024) and Zero123++ (Shi et al., 2023a) multi-view images
@@ A.5.2 Prompts for MV-LLaVA Instruct Tuning
quality estimation What do you think about the overall quality of view consistency of three synthesized novel views? Choosing from ”poor”, ”relatively poor”, ”boardline”, ”relatively good”, ”good”, ”perfect”.
Table 9: Instruct tuning prompt for Objaverse (Deitke et al., 2023) rendered multi-view images
@@
quality estimation What do you think about the overall quality of this 3D model? Choosing from ”poor”, ”relatively poor”, ”boardline”, ”relatively good”, ”good”, ”perfect”.
scale tag What do you think about the scale of the 3D model represents? Choosing from ”single_object”, ”multi-object”, ”small_scene”, ”large_scene”.
style tag What do you think about the overall style of the 3D model? Choosing from ”CAD”, ”Cartoon”, ”Photo_realistic”.
Figure 20: A test conversation with GPT-4V (OpenAI, 2023a) of evaluating generated objects.
A.6 GPT-4V based 3D Object Generation Evaluation.
We adopt method proposed in GPTeval3D (Wu et al., 2024) for more thorough and human-aligned evaluation of the quality of generated object by different methods. A full test case is shown in Fig. 20. Left 9-view image is rendered from object generated by Bootstrap3D and the right one generated by Instant3D (Li et al., 2023a). We ask GPT-4V to mainly evaluate through comparison based on three dimensions: text-image alignment, low-level texture quality and 3D plausibility. The answer of GPT-4V shows its in depth perception ability of given reasonable comparison well aligned with human preference. We thus choose to use GPT-4V rather than human volunteers to give reasonable evaluation.
We adopt the 110 test prompts proposed in GPTeval3D (Wu et al., 2024) to test Bootstrap3D generated object comparing with Instant3D (Li et al., 2023a), Zero123++ (Shi et al., 2023a) and MVDream (Shi et al., 2023b). For each methods, we conditioned model based on 110 test prompts with 4 different seeds, with each methods generates 440 objects, we make 1-to-1 comparison following aforementioned test setting. Results are reported in Tab. 10. Except MVDream (Shi et al., 2023b) (SDS) (which generates single object consuming 30 mins while Bootstrap3D only need 5 seconds.). Bootstrap3D excels in all three evaluation dimensions, which proves the ability of Bootstrap3D in creating high quality 3D objects.
Table 10: GPT-4V based evaluation result. the result is in format of ”number of objects preferred geneated by Bootstrap3D/ that of other methods”. Cases when GPT cannot answer the question or generates ”cannot decide” answer are excluded.
@@
Compared to Instant3D (Li et al., 2023a) (unofficial) 247 / 116 202 / 162 259 / 110
Compared to Zero123++ (Shi et al., 2023a) 192 / 143 210 / 161 231 / 139
Compared to MVDream (Shi et al., 2023b) (GRM) 290 / 71 245 / 131 284 / 102
Compared to MVDream (Shi et al., 2023b) (SDS) 188 / 155 173 / 190 192 / 150
A.7 Improving Direct 3D Generative Models
Figure 21: Fine tuned Shape-E generation results that shows better object-text alignment than original Shape-E (Jun & Nichol, 2023) and finetuned version in Cap3D (Luo et al., 2024).
Table 11: Test results on Shape-E. More accurate and descriptive 3D caption help model to achieve better object-text alignment.
Method FID $\downarrow$ CLIP score $\uparrow$ CLIP-R-precision $\uparrow$
Shape-E 37.2 80.4 20.3
Cap3D 35.5 79.1 20.0
Ours 35.3 81.2 22.1
In addition to fine-tuning the multiview diffusion model, we also evaluate our framework on direct 3D generative models, circumventing the use of multi-view images as intermediaries. For this purpose, we selected the Shape-E (Jun & Nichol, 2023) model for experiment and assess the outcomes following the testing method the same to Cap3D (Luo et al., 2024). Specifically, we fine-tune Shape-E using 250K BS-Objaverse data, ensuring that all entries scored greater than 3, accompanied by more precise and descriptive captions. The metrics for training and testing are consistent with those employed in Cap3D (Luo et al., 2024). Some qualitative results are presented in Fig.21, where our finetuned verson can generate object that follow text prompt more precisely. Quantitative results are detailed in Tab.LABEL:table:finetune_on_shape_e, where more accurate and desciptive captions than Cap3D can significantly improve metrics like CLIP score. Our findings indicate that improved data quality can significantly enhance object-text alignment and visual quality of Shape-E. This experiment substantiates that our pipeline, characterized by detailed captions and quality filtering, is also effective for direct 3D objects generation represented by neural field.
A.8 More Results Visualization
A.8.1 Comparison with Other Methods
@@
A.8.2 Visualization of Generated objects with Different Styles